# ViLLa: Video Reasoning Segmentation with Large Language Model

### v2

@@ Abstract
Although video perception models have made remarkable advancements in recent years, they still heavily rely on explicit text descriptions or pre-defined categories to identify target instances before executing video perception tasks. These models, however, fail to proactively comprehend and reason the user’s intentions via textual input. Even though previous works attempt to investigate solutions to incorporate reasoning with image segmentation, they fail to reason with videos due to the video’s complexity in object motion. To bridge the gap between image and video, in this work, we propose a new video segmentation task — video reasoning segmentation. The task is designed to output tracklets of segmentation masks given a complex input text query. What’s more, to promote research in this unexplored area, we construct a reasoning video segmentation benchmark. Finally, we present ViLLa: Video reasoning segmentation with Large Language Model, which incorporates the language generation capabilities of multimodal Large Language Models (LLMs) while retaining the capabilities of detecting, segmenting, and tracking multiple instances. We use a temporal-aware context aggregation module to incorporate contextual visual cues to text embeddings and propose a video-frame decoder to build temporal correlations across segmentation tokens. Remarkably, our ViLLa demonstrates capability in handling complex reasoning and referring video segmentation. Also, our model shows impressive ability in different temporal understanding benchmarks. Both quantitative and qualitative experiments show our method effectively unlocks new video reasoning segmentation capabilities for multimodal LLMs. The code and dataset will be available at https://github.com/rkzheng99/ViLLa.
@@ 1 Introduction
Figure 1: We propose a new video segmentation task that involves complex reasoning for models. Our ViLLa is an effective and efficient LLM capable of segmenting and tracking multiple objects while understanding the user’s implicit instructions.
@@
Capitalizing on the achievements of Large Language Models (LLMs) vicuna ; chatgpt ; openai2023gpt4 ; zhang2023llama , the development of large multimodal models (LMMs) Qwen-VL ; li2023blip ; liu2023llava ; 2023interngpt ; wang2023internvid ; zhu2023minigpt has notably enhanced visual perception capabilities and user-interaction experiences to new heights. However, in the absence of explicit instructions specifying target objects or categories, the majority of models alayrac2022flamingo ; li2023otter ; li2023blip ; liu2023improved ; liu2023visual ; ye2023mplug ; zhu2023minigpt , including both perception models and LMMs, fall short of generating detailed and granular responses such as instance masks, and instead produce merely general textual descriptions for images or videos. This failure, unfortunately, hinders the practical application of multimodal systems in substantial industrial and daily life scenarios, such as autonomous driving, image/video editing, robotics, and augmented reality.
Recent studies lai2023lisa ; ren2023pixellm ; yang2023improved have explored the implementation of LLMs to produce object masks in a novel reasoning segmentation task, which enhances the applicability to real-world applications. Unlike traditional segmentation, which explicitly specifies target objects (e.g., “orange”), reasoning segmentation necessitates complex reasoning to interpret more sophisticated instructions (e.g., “the fruit high in Vitamin-C”). These new approaches propose efficient LMMs that are capable of producing multiple open-set targets while maintaining diverse reasoning complexities. Albeit the success gained by these methods, they still merely focus on image-level reasoning and understanding, which fail to extend the fine-grained reasoning capabilities to the temporal dimension. The adaptation from image to video seems intuitive but still has a major obstacle: The complex relations inherent in video frames pose a challenge for image-based reasoning methods to effectively model the temporal relationships between instances using a single segmentation token or fixed segmentation codebook. Yet, concentrating on modeling temporal relationships will compromise the spatial comprehension per frame. Considering the success of using LLMs in image-level reasoning segmentation, we intuitively come up with the question: how to efficiently utilize the temporal relations across frames while maintaining the frame-level comprehension when we manage to adapt LLMs to video scenarios?
In this paper, we introduce ViLLa, an effective and efficient LMM for video reasoning segmentation and understanding which proficiently perceive and reason videos without losing the spatial details of images (As shown in Fig 1). Our core idea is to represent the videos with both temporally correlated and contextually preserved tokens. Specifically, we propose a context attention module to aggregate frame context information from visual features to text embeddings to generate contextually enriched text embeddings. To incorporate temporal relations to segmentation tokens, we design a video-frame decoder to communicate the segmentation embeddings of a single frame with the embeddings representing the whole video frame. Remarkably, ViLLa demonstrates competitive performance among all video understanding benchmarks, demonstrating the robustness of our models. As shown in  1, our model shows capability in various complicated video scenarios involving complex reasoning and multi-target tracking while segmenting. Tested on Refer-Youtube-VOS and Youtube-VIS, our model could surpass previous state-of-the-art methods by 3.6 and 2.8 in $\mathcal{J}\&\mathcal{F}$ and AP.
In addition, to validate its effectiveness, we establish a benchmark for video reasoning segmentation evaluation, called VideoReasonSeg. Comprising over one thousand video-instruction pairs, this benchmark offers persuasive evaluation metrics for the task. Utilizing a GPT-4V openai2023gpt4 -aided data curation pipeline, we generate 1934 video samples for video reasoning segmentation in either the form of multiple-choice QAs or video-instruction-mask data samples.
@@
  • We introduce the reasoning video segmentation task, which necessitates pixel-level video reasoning based on implicit user instructions. Developing such capability is a meaningful leap forward for building an omnipotent video perception model.
  • We have developed VideoReasonSeg, a comprehensive benchmark for video reasoning segmentation that comprises 1934 video-instruction-mask data samples. This benchmark plays a crucial role in evaluating and fostering the exploration of reasoning capabilities in video-based models by the research community.
  • We present our model, ViLLa, which serves as a novel LMM for video reasoning segmentation. It handles tasks with diverse reasoning complexities, achieving state-of-the-art results across various video understanding benchmarks and our proposed benchmark.
@@ 2 Related Works
Large Multimodal Models. Large multimodal models (LMMs) have made significant strides in tasks that require understanding diverse modalities. These models can be broadly categorized into two groups based on their utilization of large language models (LLMs).
@@
The first category comprises models lu2022unified ; wang2022ofa ; yu2022coca that are either trained from scratch or leverage smaller language models like BERT for text processing. These models typically employ a combination of contrastive and generative objectives to tackle a range of multimodal tasks (e.g., Coca yu2022coca ). However, their limited language understanding capacity often restricts their performance in tasks that demand massive commonsense reasoning abilities.
@@
The emergence of LLMs in the recent few years has paved the way for a new paradigm in LMM development, in which LLMs are augmented with multimodal comprehension abilities zhang2022opt ; touvron2023llama ; alpaca ; peng2023instruction . This approach typically includes utilizing adapters to align visual and textual representations within LLMs, as exemplified by models such as Flamingo alayrac2022flamingo , BLIP-2 li2023blip , MiniGPT-4 zhu2023minigpt , llama-adapter llama-adapter-v2 , LLaVA liu2023llava , InstructBLIP instructblip , InternGPT 2023interngpt , QwenVL Qwen-VL , InternVideo2 wang2024internvideo2 . Some video-related MLLM (VideoLLM) VideoChatGPT maaz2023video , and Valley luo2023valley utilize ChatGPT chatgpt to generate video instruction-tuning data, aiming to enhance instruction-following capabilities for real-world video comprehension. Even though these models have demonstrated improved performance in vision-language tasks through instructional tuning, their primary limitation lies in generating only textual outputs about the entire image or video, thereby constraining their applicability in tasks that require more fine-grained, region-level, or pixel-level understanding.
Video Segmentation. Video Instance Segmentation (VIS) yang2019video is a fundamental while challenging vision task that aims to detect, segment, and track object instances inside videos based on a set of predefined object categories simultaneously. With the prosperous video media, VIS has attracted various attention due to its numerous important applications in areas such as video understanding, video editing, autonomous driving, etc. Numerous studies athar2020stem ; lin2021video ; wu2022efficient ; wang2021end ; cheng2022masked ; heo2022vita ; wu2022seqformer ; heo2022generalized ; li2022videoknet ; OMGSeg ; dvisdaq have proposed diverse designs (such as tracking by detection, masked attention, object token association, contrastive memory, and more) to effectively model temporal relationships while classifying instances. To name a few, IDOL wu2022defense , which built upon Deformable-DETR zhu2020deformable , incorporates a contrastive learning head to obtain unique embeddings for associations he2020momentum . Subsequently, CTVIS ying2023ctvis improved IDOL by establishing a unified framework for training and inference. DVIS zhang2023dvis , on the other hand, proposed a decoupling strategy for VIS by dividing it into three independent sub-tasks: segmentation, tracking, and refinement.
@@
The Referring Video Object Segmentation (RVOS) task was introduced to segment the target instance referred by a given text in given videos. Early works khoreva2019video ; seo2020urvos used the spatial-temporal memory mechanism to strengthen temporal referring relations. The researchers then started building off-line reference models wang2019asymmetric ; wang2020context ; wu2022multi ; yan2024referred that take the whole clip as input. Currently, query-based Transformer models botach2022end ; wu2022language have begun to dominate. MTTR botach2022end utilized an instance-level segmentation pipeline to predict sequences for all instances and determine the most suitable sequence corresponding to the referred object. ReferFormer wu2022language transformed text expressions into queries to attend to relevant regions in videos.
@@
Reasoning Segmentation. In numerous practical applications, it is essential to comprehend visual inputs at a more fine-grained level, such as specific regions or even individual pixels. To address this need, several approaches have been proposed to equip Large Language Models (LLMs) with fine-grained understanding capabilities. Kosmos-2 peng2023kosmos , InternGPT 2023interngpt , and Ferret you2023ferret offer grounding capabilities to specified image regions, typically encoding location coordinates as tokens for integration with LLMs. Yet, these approaches still fail to provide pixel-level outputs. To bridge this gap, LISA lai2023lisa integrates SAM kirillov2023segment with LLMs to tackle segmentation tasks. Furthermore, LISA pioneered the use of Large Multimodal Models (LMMs) for complex instruction reasoning, a paradigm shift from traditional tasks that rely on explicit human instructions for object or category identification. This innovative approach enables LMMs to reason about intricate instructions, moving beyond simplistic object recognition. In order to generate masks for reasoning tasks involving multiple targets, PixelLM ren2023pixellm proposes a novel pixel decoder and a comprehensive segmentation codebook to conduct multi-target reasoning segmentation. However, these methods mainly focus on reasoning image segmentation and fail to reason and deduct at a pixel-level across temporal dimensions. When they are transferred to the video domain, the segmentation token is not capable of modeling multiple targets as well as their moving trajectories. To bridge this gap, we propose ViLLa with context aggregation and video-frame decoder to address the problem of video reasoning segmentation.
3 Video Reasoning Segmentation
3.1 Problem Definition
Video reasoning segmentation task aims to output a binary segmentation mask sequence $\{m_{t}\}_{t=1}^{T}\in\mathbb{R}^{T\times H\times W}$, given an input video clip consisting of $T$ frames $V\in\mathbb{R}^{T\times H\times W\times 3}$ and an implicit query text instruction $\mathbf{x}_{txt}$. The task shares a similar formulation with the referring video object segmentation, but there’s a major difference that makes video reasoning segmentation much more challenging: The complexity of the input query text. Apart from the straightforward text description, such as “a zebra to the left”, the query text of video reasoning segmentation includes more complex and nuanced descriptions of the instance (e.g., “a type of African mammal known for their distinctive black and white striped coat patterns walking at the end of its herd”), which involves complicated reasoning as well as the world knowledge. Also different from the image reasoning segmentation, the query text of video reasoning segmentation includes more motion information, such as “a person wearing a white shirt is in the ocean on a surfboard riding a wave”, which requires the model to capture the moving trajectory of the “person riding a wave”. Thus, this topic is challenging and worth exploring, especially in a time when video inputs have become ubiquitous.
3.2 Benchmark
Given the lack of quantitative evaluation, it is necessary to establish a benchmark for the video reasoning segmentation task. To ensure reliable assessment, we have collected a diverse set of videos from Youtube-VIS yang2019video , OVIS qi2022occluded , and LV-VIS wang2023LVVIS , annotating them with implicit text instructions and high-quality target masks. To fully evaluate the reasoning capabilities, we design two types of evaluation standards: 1) multiple-choice QAs; 2) instructions and answers. Considering that not all the annotations of selected datasets follow the multiple-choice QA format, we automatically convert the video annotations into this format via LLMs. This is mainly because the open-ended answer has to be scored by LLMs or user studies, which may either introduce evaluation bias or manual intervention. Ultimately, we produce 2 multiple-choice QA pairs for each of the video clips. As for the answers, we mainly focus on the quality of segmentation masks. The resulting VideoReasonSeg benchmark comprises a total of 1934 video-instruction-mask data samples. This dataset is further partitioned into three splits: train, val, and test, containing 1000, 400, and 534 video samples, respectively. The details of data annotation are given in the supplementary material.
Dataset Generation Pipeline. Inspired by PixelLM, we use GPT-4V and Qwen-VL for generating questions and multiple choices, because of their superiority in understanding visual contents. These models have been instrumental in generating more nuanced, contextual, and naturalistic questions. Specifically, we feed all the instance category names and corresponding masks in the video information in the video to GPT-4V. With our carefully designed prompts, GPT-4V autonomously selects instances to construct question-answer pairs based on the video contents. Examples of such prompts are illustrated in the supplementary files.
4 Method
@@
Figure 2: Overall framework of the proposed ViLLa. Given the input video frames and user input text query, the context attention module firstly aggregates text-related visual cues from the visual features to the text embeddings. The large language model generates text output and segmentation tokens with the input of visual features, visually enriched text embeddings, and pre-defined segmentation tokens. Finally, the segmentation tokens are fed to the Video-Frame Decoder to produce the final output segmentation mask tracklets.
@@ 4.1 Model Design
Model Overview. As depicted in Fig. 2, ViLLa features a streamlined architecture, comprising mainly several main parts: 1) a pre-trained visual encoder that aligns with a text encoder 2) a large language model 3) a context attention module and 4) a video-frame interaction decoder.
@@
While 1) and 2) components follow well-established large language model (LLM) architectures to ensure compatibility, the context aggregation module (CAM) and the Video-Frame decoder are crucial in equipping the LLMs with the capacity to generate masks across temporal dimensions. We utilize the context aggregation module to aggregate the target-text-relevant visual features of the current frame to the input user text embeddings. The output text embeddings that carry the most crucial visual cues are then fed into the large language models along with the visual features. In video-frame decoder, we use the video-scale segmentation tokens to interact with frame-level tokens to produce the final refined complete segmentation tokens that contain both video-level and frame-level information. Finally, the segmentation tokens produce the prediction masks in conjunction with the multi-scale video features.
@@
Encoders. For input frame $V_{\text{t}}$, the visual encoder $\mathcal{I}$ extracts multi-scale visual features $F_{{t}}=\{F_{{t}}^{\ell}\}_{\ell=1}^{L}$ from $\mathcal{I}(V_{{t}})$, comprising $L$ visual features output at select layers of $\mathcal{I}$. Meanwhile, the user instruction is fed into the text encoder with the visual embedding $F_{t}^{L}\in\mathbb{R}^{N\times C}$ (where $N=H/p\times W/p$ and $C$ indicate the number of frame patches and embedding channels, respectively) and generates the $\mathbf{x}_{txt}\in\mathbb{R}^{M\times C}$, where $M$ denotes the number of queries. In this part, we choose QFormer as the text encoder to maintain the cross-modality alignment and interaction. By incorporating this approach, the text embeddings now capture visual cues that are highly relevant to the user’s instructions. The output of the final layer, $F_{t}^{L}$, encodes global image information and is transformed to align with the language space of LLMs via a vision-to-language projection layer $\text{Proj}_{V\rightarrow L}$.

### v3

@@ Abstract
Recent efforts in video reasoning segmentation (VRS) integrate large language models (LLMs) with perception models to localize and track objects via textual instructions, achieving barely satisfactory results in simple scenarios. However, they struggled to discriminate and deduce the objects from user queries in more real-world scenes featured by long durations, multiple objects, rapid motion, and heavy occlusions. In this work, we analyze the underlying causes of these limitations, and present ViLLa: Video reasoning segmentation with Large Language Model. Remarkably, our ViLLa manages to tackle these challenges through multiple core innovations: (1) a context synthesizer that dynamically encodes the user intent with video contexts for accurate reasoning, resolving ambiguities in complex queries, and (2) a hierarchical temporal synchronizer that disentangles multi-object interactions across complex temporal scenarios by modelling multi-object interactions at local and global temporal scales. To enable efficient processing of long videos, ViLLa incorporates (3) a key segment sampler that adaptively partitions long videos into shorter but semantically dense segments for less redundancy. What’s more, to promote research in this unexplored area, we construct a VRS benchmark, VideoReasonSeg, featuring different complex scenarios. Our model also exhibits impressive state-of-the-art results on VideoReasonSeg, Ref-YouTube-VOS, Ref-DAVIS17, MeViS, and ReVOS. Both quantitative and qualitative experiments demonstrate that our method effectively enhances video reasoning segmentation capabilities for multimodal LLMs.
@@ 1 Introduction
Capitalizing on the achievements of Large Language Models (LLMs) [11, 39, 40, 76], the development of large multimodal models (LMMs) [3, 26, 33, 35, 59, 81] has notably enhanced visual perception capabilities and user-interaction experiences to new heights. However, in the absence of explicit instructions specifying target objects or categories, the majority of models [1, 25, 26, 32, 33, 72, 81], including both perception models and language ones, struggle to generate detailed and granular outputs such as instance masks. Instead, these models typically generate only general textual descriptions for images or videos. This limitation hinders the practical application of multimodal systems in various scenarios, such as autonomous driving [58], image and video editing [31], robotics [49], and augmented reality [47].
Recent studies [24, 46, 71] have investigated the application of LLMs to generate object masks in a novel reasoning segmentation task that requires complex reasoning to interpret sophisticated instructions. These approaches introduce efficient language models capable of producing multiple open-set targets while accommodating diverse reasoning complexities. However, despite their successes, these methods remain limited to image-level reasoning and do not extend fine-grained reasoning capabilities to the temporal dimension. To address this issue, researchers [68, 4] managed to equip the image reasoning segmentation model with temporal understanding ability, and propose VRS. VISA [68] employs a video understanding model [29] to select the key frame for LISA [24] to segment and subsequently utilizes an object tracker [9] to maintain temporal consistency. However, the quality of segmentation is highly dependent on the accuracy of key frame selection, which is susceptible to error accumulation through propagation as the number of input frames increases. On the other hand, VideoLISA [4] adopts a sparse dense sampling strategy to compress visual features and extends the single segmentation token to track an instance across frames, especially when the movement is rapid. Nevertheless, the sampling strategy fails to adaptively select the key video segments, making it challenging to capture the essential information from the video. Also, because a single segmentation token is unable to segment and track multiple instances across frames, the segmentation and tracking quality decrease significantly when the video scene becomes intricate (e.g., involving multiple objects with occlusion), as proven in various previous Video Instance Segmentation (VIS) approaches [21, 79], which utilize multiple-level of tokens to segment and track instances.
(a) Refer to caption
(b) Refer to caption
(c) Refer to caption
Figure 1: Our ViLLa is an effective and efficient LMM capable of segmenting and tracking: (a) multiple objects with rapid motion; (b) objects in crowded scenes; (c) objects in long videos with occlusions.
To tackle the aforementioned challenges, we propose ViLLa, an effective and efficient LMM for video reasoning segmentation. ViLLa proficiently perceives and reasons videos with multiple objects, rapid motion, and diverse reasoning complexities (as shown in Fig. 1). Specifically, our ViLLa has three key components: Key Segment Extractor, Context Synthesizer, and Hierarchical Temporal Synchronizer. These designs put effort into solving the complex video reasoning segmentation from different perspectives. The Key Segment Extractor selects the most relevant video segments, which reduces redundant visual tokens and effectively turns long videos into shorter segments. To make the responses more pertinent to users’ intentions, Context Synthesizer synthesizes frame context information from visual features to text embeddings to generate contextually enriched text and highly input-responsive visual embeddings. To build up stronger temporal relations to multi-target segmentation tokens in complex scenes, the Hierarchical Temporal Synchronizer communicates and consolidates the multi-scale segmentation embeddings across temporal dimension via attention-based fusion strategy. Tested on multiple referring video object segmentation benchmarks, our model could surpass previous state-of-the-art LLM-based methods by a large margin. In addition, to validate its effectiveness, we establish a benchmark for video reasoning segmentation evaluation, called VideoReasonSeg. Utilizing a GPT-4V [40]-aided data curation pipeline, we curate 3k high quality video samples for video reasoning segmentation spanning multiple-choice QAs and video-instruction-mask samples, and the video scenarios contain diverse reasoning complexities (as presented in the appendix), which offers convincing metrics.
@@
  • We analyze the limitations of current VRS approaches and propose a framework named ViLLa that segments and tracks the desired objects in more complex scenarios.
  • We develop three key modules that resolves complex scenario reasoning from different perspectives, namely key segment extractor, context synthesizer, and hierarchical temporal synchronizer. Our extractor selects the most relevant video segments from the complex input video, while the synthesizer combines and infuses user intention and context information, and the synchronizer communicates multi-scale segmentation tokens spatio-temporally.
  • We develop VideoReasonSeg, a comprehensive benchmark for VRS that spans different complex reasoning scenarios, comprising 3k videos and 15k object descriptions for instruction tuning and evaluation. This benchmark plays a crucial role in evaluating and fostering the exploration of reasoning capabilities in video-based models.
  • We conduct extensive experiments on challenging referring segmentation benchmarks, ReVOS, and our proposed VideoReasonSeg. The state-of-the-art results demonstrate the effectiveness of our method.
@@ 2 Related Works
Large Multimodal Models. Large multimodal models (LMMs) can be broadly categorized into two groups based on their utilization of large language models (LLMs). The first category comprises models [36, 57, 75] that are either trained from scratch or leverage smaller language models like BERT for text processing. These models typically employ a combination of contrastive and generative objectives to tackle a range of multimodal tasks (e.g., Coca [75]). However, their limited language understanding capacity often restricts their performance in tasks that demand massive commonsense reasoning abilities.
@@
The emergence of LLMs in the recent few years has paved the way for a new paradigm in LMM development, in which LLMs are augmented with multimodal comprehension abilities [77, 52, 51, 41]. This approach typically includes utilizing adapters to align visual and textual representations within LLMs, as exemplified by models such as Flamingo [1], BLIP-2 [26], MiniGPT-4 [81], LLaMA-Adapter [17], LLaVA [33], InstructBLIP [12], InternGPT [35], QwenVL [3], InternVideo2 [60]. Some video-related MLLM (VideoLLM) VideoChatGPT [38], and Valley [37] utilize ChatGPT [39] to generate video instruction-tuning data, aiming to enhance instruction-following capabilities for real-world video comprehension. Even though these models have demonstrated improved performance in vision-language tasks through instructional tuning, their primary limitation lies in generating only textual outputs.
@@
Video Segmentation. Video Instance Segmentation (VIS) [70] aims to detect, segment, and track object instances inside videos based on a set of predefined object categories simultaneously. Numerous studies [30, 64, 61, 27, 80] have proposed diverse designs (such as tracking by detection [2], masked attention [8], object token association [66, 21, 28], contrastive memory [20, 73], and more) to effectively model temporal relationships. For example, IDOL [67], which is built upon Deformable-DETR [82], incorporates a contrastive learning head to obtain unique embeddings for associations [18]. DVIS [79], on the other hand, proposed a decoupling strategy by dividing VIS into image segmentation, tracking, and refinement. The Referring Video Object Segmentation (RVOS) task was introduced to segment the target instance referred by a given text in given videos. Early works [22, 48] used the spatial-temporal memory mechanism to strengthen temporal referring relations. The researchers then started building offline reference models [55, 54, 62, 69] that take the whole clip as input. Currently, query-based Transformer models [5, 65] begin to dominate. MTTR [5] utilized an instance-level segmentation pipeline to determine the most suitable sequence corresponding to the referred object. ReferFormer [65] transformed text expressions into queries to attend to relevant regions in videos.
@@
Reasoning Segmentation. In numerous practical applications, it is essential to comprehend visual inputs at a more fine-grained level, such as specific regions. Kosmos-2 [42], InternGPT [35], and Ferret [74] offer grounding capabilities to specified image regions. Yet, these approaches still fail to provide pixel-level outputs. To bridge this gap, LISA [24] integrates SAM [23] with LLMs to tackle segmentation tasks. PixelLM [46] proposes a novel pixel decoder and a segmentation codebook to generate masks for multiple targets. Then, GLaMM [45] proposes region-level caption and segmentation tasks. Several works [15, 34, 44] use LLMs as agents to assign different visual experts. OMG-LLaVA [78] implements a new comprehensive baseline that has only a single visual encoder. However, these methods mainly focus on reasoning image segmentation and fail to reason across temporal dimensions. When they are transferred to the video domain, the segmentation token is not capable of modeling multiple targets as well as their moving trajectories. Early approaches, namely VISA [68] and VideoLISA [16] manage to bridge the gap by implementing different frame sampling strategies, but VISA suffers from the error accumulation from key frame selection and VideoLISA fails to adaptively extract relevant frames which leads to redundancy. After analyzing their drawbacks, we propose ViLLa with key segment extractor, context synthesizer, and hierarchical temporal synchronizer to address the problem.
@@
Figure 2: Overall framework of ViLLa. Given the input video frames and user input text query, the Key Segment Extractor is proposed to select the most query-relevant video segments from the video. Then, the Context Synthesizer aggregates text-related visual cues from the visual features to the text embeddings and selects the relevant visual features. The large language model generates text output and segmentation tokens with the input of visual features, visually enriched text embeddings, and pre-defined segmentation tokens. Finally, the segmentation tokens are fed to the Hierarchical Temporal Synchronizer to produce the final output segmentation mask tracklets.
3 ViLLa
Video reasoning segmentation aims to output a binary segmentation mask sequence $\{m_{t}\}_{t=1}^{T}\in\mathbb{R}^{T\times H\times W}$, given an input video clip consisting of $T$ frames $V\in\mathbb{R}^{T\times H\times W\times 3}$ and a query instruction $\mathbf{x}_{txt}$. The task shares a similar formulation with the referring video object segmentation (VOS), but it has complex queries that require reasoning.
@@ 3.1 Model Design
Model Overview. As depicted in Fig. 2, ViLLa features a streamlined architecture, comprising mainly several main parts: 1) a pre-trained visual encoder that aligns with a text encoder, 2) a large language model (LLM), 3) a key segment extractor, 4) a context synthesizer, and 5) a hierarchical temporal synchronizer.
@@
The key segment extractor aims to extract the pivotal video segments from a long video input. The context synthesizer and the hierarchical temporal synchronizer are crucial in equipping the LLMs with the capacity to generate masks across temporal dimensions. We utilize the context synthesizer to aggregate the target-text-relevant visual features of the current frame to the input text embeddings. The output text embeddings that carry the most crucial visual cues are then fed into the LLM along with the visual features. In the hierarchical temporal synchronizer, we use the video-scale segmentation tokens to interact with frame-level tokens to produce the final refined complete segmentation tokens that contain both video-level and frame-level information. Finally, the segmentation tokens produce the prediction masks in conjunction with the multi-scale video features.
@@
Encoder. For input frame $V_{\text{t}}$, the visual encoder $\mathcal{I}$ extracts multi-scale visual features $F_{{t}}=\{F_{{t}}^{\ell}\}_{\ell=1}^{L}$ from $\mathcal{I}(V_{{t}})$, comprising $L$ visual features at the selected layers of $\mathcal{I}$. Meanwhile, the user instruction is fed into the text encoder with the visual embedding $F_{t}^{L}\in\mathbb{R}^{N\times C}$ (where $N=H/p\times W/p$ and $C$ indicate the number of frame patches and embedding channels, respectively) and generates the $\mathbf{x}_{txt}\in\mathbb{R}^{M\times C}$, where $M$ denotes the number of queries. We choose QFormer as the text encoder to maintain the cross-modality alignment. By incorporating this approach, the text embeddings now capture visual cues that are highly relevant to the user’s instructions. The output of the final layer, $F_{t}^{L}$, encodes global image information and is transformed to align with the language space of LLMs via a vision-to-language projection layer $\text{Proj}_{V\rightarrow L}$.
3.2 Key Segment Extractor
For short videos that typically has very few actions and objects, the segmentation can be solved by image-level reasoning segmentation methods in a frame-by-frame style. These cases, however, can’t reflect the temporal reasoning capability of video reasoning segmentation models. We argue that the capability of handling complex videos, which include occluded scenes and longer duration, truly reflects the model’s temporal reasoning capacity. As shown in Fig. 2, the user input could be answered by segments that contain key information, while the rest of the video is irrelevant. Stemming from this idea, we implement Grounded-VideoLLM [53], a SOTA multi-modal LLM capable of grounding the input user queries in long videos, to serve as our Key Segment Extractor. We use a prompt specifying the user query and ask the model to output the starting and ending timestamps of the segments that describe the content of the user query. We take the average from the top-K responses to obtain the starting and ending frames of the key segments, where we denote as $V_{key}$, comprising $T_{key}$ frames. Based on the key segments, we also sample $T_{ref}$ using an adaptive global sampling strategy. We will specify the details of different sampling strategies in the appendix. The extracted frames are then fed into the encoder.
3.3 Context Synthesizer